# Cameras as Relative Positional Encoding

### v1

@@ 5 Conclusion and Future Work
In this work, we presented a study on camera conditioning for multiview transformers, with a focus on relative encodings that model relationships between viewpoints. Our experiments highlight how representing cameras as relative positional encodings—particularly relative encodings that capture both camera intrinsics and extrinsics—improves multiview transformer performance and generalization characteristics across settings and tasks. We hope that these results can help guide the design of future multiview computer vision models.
Possibilities for future work include both applications of relative camera conditioning and investigation of conditioning methods like PRoPE. Possible questions include numerical stability when directly multiplying projective matrices with Q/K/V vectors; it may be possible, for example, for ill-conditioned matrices to emerge from telephoto focal lengths. Furthermore, incorporating “multi-frequency” [19] encoding for camera parameters remains nontrivial due to the non-commutativity of projective transforms, a challenge shared by relative SE(3) attention methods [3, 4].
@@ A.1.1 LVSM-based Model Details
We adhere to the original LVSM [8] implementation specifications, maintaining consistency across most settings. For complete details regarding these configurations, we direct readers to the original LVSM paper [8]. The modifications we made include:
@@ A.1.2 UniMatch Modification Details
UniMatch [14]’s model consists of a cross-attention transformer to capture multi-view relationship, as well as a self-attention transformer that serves as a single-view image encoder/decoder. We inject the camera information into both transformers, on the Q/K/V/O vectors, using our formulation. Note that our formula exactly falls back to RoPE [28] in the single-view scenario. It therefore naturally works with both transformer networks in UniMatch.
@@ A.1.3 Spatial Cognition Model Details
We formulate this task as a classification problem, where the number of classes corresponds to the number of input image-camera pairs, and the “ground-truth” class is the ID of the inconsistent image-camera pair. The training objective is to identify the inconsistent pair, which we optimize using the cross-entropy loss. The architecture of the model is largely similar to LVSM [8], where the only change is that the last linear layer is modified to output a single scalar for each token. Outputs are then averaged over all tokens for each input pair, resulting in a vector that represents the probability that each input pair is the “bad” one (over softmax). We provide more data exemplars we used for training and testing in Figure A.3.
@@ A.2.1 Ablating PRoPE
As our proposed PRoPE includes two terms: the projective relationship between cameras ($\mathbf{D}_{t}^{\text{Proj}}$) and the patch coordinate relationship ($\mathbf{D}_{t}^{\text{RoPE}}$). We here ablate each of them to study their contribution. As shown in Table A.1, as a crucial component to the system, modeling the projective relationship between cameras alone already yields strong performance. Introducing RoPE [28] further enhances model’s ability on understanding patch relationship, which is a minimal unit in vision transformer architecture. Notably, in PRoPE, we allocate half of the feature channels to encode each term. Thus, ablating one term means using all feature channels to encode the remaining one – result a fair comparison with constant number of model parameters.
@@
Table A.1: Ablation Study on PRoPE. $\mathbf{D}_{t}^{\text{Proj}}$ is crucial for encoding the relative camera information, and $\mathbf{D}_{t}^{\text{RoPE}}$ is also helpful to capture the the relative patch coordinate. Experiments are conducted on RealEstate10K [12] with CamRay as input.
A.2.2 More Results for Stereo Depth Estimation
In Figure A.2 we show more qualitative results on the task of stereo depth estimation with Uni-Match [14].
Figure A.1: More Qualitative Results of Novel View Synthesis on RealEstate10K [12] and Objaverse [13].
Figure A.2: More Qualitative Results of Stereo Depth Estimation on RGBD [61], SUN3D [62] and Scenes11 [63].

### v2

@@ 5 Conclusion and Future Work
In this work, we highlight how representing cameras as relative positional encodings—particularly in a way that captures both camera intrinsics and extrinsics—improves multiview transformers across settings and tasks. Possibilities for future work include improving numerical stability when directly multiplying projective matrices with Q/K/V vectors; it may be possible, for example, for ill-conditioned matrices to emerge from telephoto focal lengths. It would also be interesting to extend PRoPE to distorted camera models, for example by applying PRoPE to per-patch projective approximations. Furthermore, incorporating “multi-frequency” [schenck2025learning, su2024roformer] encoding for camera parameters remains nontrivial due to the non-commutativity of projective transforms, a challenge shared by relative SE(3) attention methods [miyato2023gta, kong2024eschernet].
@@ A.1.1 LVSM-based Model Details
We adhere to the original LVSM [jin2024lvsm] implementation specifications, maintaining consistency across most settings. For complete details regarding these configurations, we direct readers to the original LVSM paper [jin2024lvsm]. The modifications we made include:
@@ A.1.2 UniMatch Modification Details
UniMatch [xu2023unifying]’s model consists of a cross-attention transformer to capture multi-view relationship, as well as a self-attention transformer that serves as a single-view image encoder/decoder. We inject the camera information into both transformers, on the Q/K/V/O vectors, using our formulation. Note that our formula exactly falls back to RoPE [heo2024rotary] in the single-view scenario. It therefore naturally works with both transformer networks in UniMatch.
@@ A.1.3 Spatial Cognition Model Details
We formulate this task as a classification problem, where the number of classes corresponds to the number of input image-camera pairs, and the “ground-truth” class is the ID of the inconsistent image-camera pair. The training objective is to identify the inconsistent pair, which we optimize using the cross-entropy loss. The architecture of the model is largely similar to LVSM [jin2024lvsm], where the only change is that the last linear layer is modified to output a single scalar for each token. Outputs are then averaged over all tokens for each input pair, resulting in a vector that represents the probability that each input pair is the “bad” one (over softmax). We provide more data exemplars we used for training and testing in Figure A.3.
@@ A.2.1 Ablating PRoPE
As our proposed PRoPE includes two terms: the projective relationship between cameras ($\mathbf{D}_{t}^{\text{Proj}}$) and the patch coordinate relationship ($\mathbf{D}_{t}^{\text{RoPE}}$). We here ablate each of them to study their contribution. As shown in Table A.1, as a crucial component to the system, modeling the projective relationship between cameras alone already yields strong performance. Introducing RoPE [heo2024rotary] further enhances model’s ability on understanding patch relationship, which is a minimal unit in vision transformer architecture. Notably, in PRoPE, we allocate half of the feature channels to encode each term. Ablating one term therefore means using all feature channels to encode the remaining one.
@@
Table A.1: Ablation Study on PRoPE. $\mathbf{D}_{t}^{\text{Proj}}$ is crucial for encoding the relative camera information, and $\mathbf{D}_{t}^{\text{RoPE}}$ is also helpful to capture the relative patch coordinate. Experiments are conducted on RealEstate10K [zhou2018stereo] with CamRay as input.
A.2.2 PRoPE v.s. GTA: More Analysis on the Affect of Intrinsic Modeling
The main difference between our proposed PRoPE and prior work GTA [miyato2023gta] is the introduce of camera intrinsic into the formulation. While GTA itself does not take into account camera intrinsics, as we have pointed out in Section 4.4, it is compatible with our proposed CamRay – which encodes the intrinsic at token level – to form a complete representation for camera conditioning. In this section, we run more thorough experiments focusing comparing PRoPE and GTA in differnet settings (with/without CamRay) as the camera conditioning techniques on both UniMatch for stereo depth estimation in Table A.4 and LVSM for novel view synthesis in Table A.3, both with augmented intrinsics in training to highlight the benefits of PRoPE on intrinsic modeling.
The new results corroborate the benefits of PRoPE that were originally shown in the main paper. We find that PRoPE consistently outperforms GTA. Importantly, both PRoPE and PRoPE+CamRay outperform GTA+CamRay, despite the fact that these methods include the same information and are all invariant to the choice of global frame. This supports our finding that attention-level intrinsic conditioning (as in PRoPE) is important.
3-view 6-view 9-view
Zip-NeRF [barron2023zip] 12.77 0.271 0.705 13.61 0.284 0.663 14.30 0.312 0.633
ZeroNeRF [sargent2023zeronvs] 14.44 0.316 0.680 15.51 0.337 0.663 15.99 0.350 0.655
ReconFusion [wu2024reconfusion] 15.50 0.358 0.585 16.93 0.401 0.544 18.19 0.432 0.511
CAT3D [gao2024cat3d] 16.62 0.377 0.515 17.72 0.425 0.482 18.67 0.460 0.460
CAT3D [gao2024cat3d] + PRoPE 16.93 0.382 0.505 18.01 0.443 0.479 18.98 0.474 0.461
Table A.2: Incorporating PRoPE into CAT3D. Adding PRoPE to the CAT3D [gao2024cat3d] multiview diffusion model yields improvements over the original model, with zero additional model parameters and negligible computational overhead.
[…]
Plücker 19.89 21.03 20.75 20.32
GTA 15.77 18.14 18.12 18.29
GTA+CamRay 21.41 22.70 22.35 21.77
PRoPE 21.42 22.95 22.75 22.40
PRoPE+CamRay 21.86 23.06 22.83 22.40
PSNR ($\uparrow$) across zoom-in levels.
[…]
Plücker 0.608 0.694 0.732 0.755
GTA 0.512 0.626 0.683 0.721
GTA+CamRay 0.673 0.741 0.764 0.780
PRoPE 0.678 0.748 0.775 0.794
PRoPE+CamRay 0.693 0.751 0.776 0.794
SSIM ($\uparrow$) across zoom-in levels.
[…]
Plücker 0.327 0.272 0.308 0.345
GTA 0.641 0.488 0.439 0.403
GTA+CamRay 0.238 0.191 0.242 0.288
PRoPE 0.247 0.193 0.227 0.244
PRoPE+CamRay 0.218 0.183 0.220 0.242
LPIPS ($\downarrow$) across zoom-in levels.
[…]
UniMatch 0.234 0.228 0.238 0.299
+Plucker -2.47% -9.35% -7.74% -4.67%
+GTA -28.83% -33.48% -32.15% -12.55%
+GTA+CamRay -29.90% -35.25% -32.42% -13.65%
+PRoPE -33.75% -36.66% -33.85% -26.40%
+PRoPE+CamRay -31.99% -35.82% -33.24% -28.21%
abs_rel ($\downarrow$) across zoom-in levels.
[…]
UniMatch 0.357 0.376 0.408 0.613
+Plucker -1.08% -9.46% -2.76% -3.37%
+GTA -31.77% -41.52% -41.16% -11.88%
+GTA+CamRay -33.54% -44.06% -42.34% -13.09%
+PRoPE -45.13% -46.79% -44.53% -38.56%
+PRoPE+CamRay -41.79% -46.75% -44.08% -39.88%
sq_rel ($\downarrow$) across zoom-in levels.
[…]
UniMatch 1.059 1.031 1.035 1.188
+Plucker +1.03% -3.95% -0.53% +1.64%
+GTA -19.55% -26.12% -25.68% -4.50%
+GTA+CamRay -20.41% -27.44% -25.72% -6.12%
+PRoPE -28.64% -29.79% -28.25% -22.61%
+PRoPE+CamRay -26.86% -29.29% -28.09% -23.74%
rmse ($\downarrow$) across zoom-in levels.
[…]
UniMatch 0.322 0.322 0.336 0.421
+Plucker +0.49% -5.46% -2.33% +2.00%
+GTA -23.29% -30.33% -30.03% -3.24%
+GTA+CamRay -24.08% -31.64% -30.23% -4.34%
+PRoPE -32.01% -33.56% -32.67% -26.78%
+PRoPE+CamRay -29.98% -32.68% -32.31% -27.71%
rmse_log ($\downarrow$) across zoom-in levels.
Table A.4: Additional Stereo Depth Estimation Results on Out-of-distribution Intrinsics at Test Time. Experiments are conducted with UniMatch [xu2023unifying]’s official code as described in Section 4.6 but with 1/8 of the training resources (2 GPUs x 50k steps).
A.2.3 Qualitative Results for Stereo Depth Estimation
In Figure A.2 we show more qualitative results on the task of stereo depth estimation with Uni-Match [xu2023unifying].
Figure A.1: More Qualitative Results of Novel View Synthesis on RealEstate10K [zhou2018stereo] and Objaverse [deitke2023objaverse].
Figure A.2: More Qualitative Results of Stereo Depth Estimation on RGBD [sturm2012benchmark], SUN3D [xiao2013sun3d] and Scenes11 [ummenhofer2017demon].